# Sample Complexity of Distributionally Robust Average-Reward Reinforcement Learning

### v2

@@ 1 Introduction
Reinforcement learning (RL) (Sutton and Barto,, 2018) is a core machine learning framework in which agents learn to make decisions by interacting with their environments to maximize long-term rewards. RL has been successfully applied across a wide range of domains—from classic applications in robotics and control systems (Kober et al.,, 2013; Gu et al.,, 2017) to more recent advances in game playing (Li et al.,, 2009; Choi et al.,, 2009; Deng et al.,, 2017) and large language model (LLM)-driven reasoning tasks (Wei et al.,, 2022; Guo et al.,, 2025).
@@
To address this challenge, Zhou et al., (2021) built upon the distributionally robust Markov decision process (DR-MDP) framework (Iyengar,, 2005; Nilim and El Ghaoui,, 2005; Wiesemann et al.,, 2013) to propose a distributionally robust reinforcement learning (DR-RL) framework. Subsequent work advanced the field, including both model-free (Liu et al., 2022a, ; Wang et al., 2023b, ; Wang et al., 2024c, ) and model-based settings (Panaganti and Kalathil,, 2022; Xu et al.,, 2023; Clavier et al.,, 2024; Shi et al.,, 2024), as well as approaches for offline learning (Shi and Chi,, 2024) and generative models (Wang et al., 2023b, ; Wang et al., 2024c, ; Yang et al.,, 2022; Clavier et al.,, 2024), along with functional approximations (Blanchet et al.,, 2023; Ma et al.,, 2022).
@@
Average-reward RL is not only important but also more challenging in terms of algorithm design and theoretical analysis. In the standard (non-robust) RL setting, the minimax sample complexity for generative models in discounted-reward cases was resolved as early as 2013 Azar et al., (2013). In contrast, analogous results for average-reward settings were only developed much later, with recent advances in Wang et al., 2023a ; Zurek and Chen, (2023); Zurek and Chen, 2024a ; Zurek and Chen, 2024b ; Zurek and Chen, (2025) under granular structural assumptions.
@@ 2 Literature Review
Sample Complexity of Discounted-Reward Tabular RL: There is an extensive literature on the sample complexity of tabular reinforcement learning. In standard (non-robust) settings, the minimax sample complexity for discounted-reward problems has been well studied. Early works by Azar et al., (2013); Li et al., (2020) established the minimax rate of $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|(1-\gamma)^{-3}\varepsilon^{-2})$. More recent research has shifted toward instance-dependent bounds that leverage structural properties of the MDP. For example, Wang et al., 2024a ; Zurek and Chen, 2024b derive tighter instance-dependent bounds of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}(1-\gamma)^{-2}\varepsilon^{-2})$ and $\widetilde{O}(|\mathbf{S}||\mathbf{A}|\mathrm{H}(1-\gamma)^{-2}\varepsilon^{-2})$ under the assumptions that $P$ is uniformly ergodic or weakly communicating, where $\mathrm{H}$ denotes the span of the relative value function.
Sample Complexity of Average-Reward Tabular RL: Recently, there has been growing interest in the sample complexity of average-reward reinforcement learning in standard (non-robust) settings. Early work by Jin and Sidford, (2020) established a bound of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}^{2}\varepsilon^{-2})$ using primal-dual stochastic mirror descent. A rate of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}\varepsilon^{-3})$ was established later via a reduction to the discounted MDP setting (Jin and Sidford,, 2021). Subsequent analyses achieved tighter bounds: Wang et al., 2024a obtained a rate of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}\varepsilon^{-2})$, while Zurek and Chen, (2023); Zurek and Chen, 2024b derived instance-dependent optimal rates of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|\mathrm{H}\varepsilon^{-2})$ for weakly communicating MDPs. Further refinements (Zurek and Chen, 2024a, ; Zurek and Chen,, 2025) achieved similar rates without requiring prior knowledge, using plug-in and span penalization approaches. Beyond the generative model setting, Zhang et al., (2021) and Chen, (2025) provided finite-sample analyses for synchronous and asynchronous Q-learning, respectively. Asymptotic properties were also studied in Yu et al., (2025); Wan et al., (2024) for asynchronous $Q$-learning.
DR-DMDP and DR-RL: Our work builds on the theoretical foundations of robust MDPs (González-Trejo et al.,, 2002; Iyengar,, 2005; Nilim and El Ghaoui,, 2005; Wiesemann et al.,, 2013; Xu and Mannor,, 2010; Shapiro,, 2022; Wang et al., 2024b, ), which primarily develop dynamic programming principles under the discounted-reward setting. Recent advances in distributionally robust reinforcement learning (DR-RL) have investigated the sample complexity of DR-DMDPs under various divergence-based uncertainty sets. For example, Wang et al., 2024c ; Shi and Chi, (2024) establish a model-free upper bound of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|(1-\gamma)^{-4}\varepsilon^{-2})$ under KL-divergence. Under $\chi^{2}$-divergence, Shi et al., (2024) obtain a similar upper bound, while Clavier et al., (2024) show that $l_{p}$-norm constraints admit a tighter minimax rate of $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|(1-\gamma)^{-3}\varepsilon^{-2})$. Several other works contribute to the theoretical and algorithmic landscape of DR-RL in various settings, including Panaganti and Kalathil, (2021); Yang et al., (2021); Xu et al., (2023); Blanchet et al., (2024); Liu et al., 2022b ; Wang et al., 2023c ; Yang et al., (2023).
Distributionally Robust Average-Reward MDPs: While the sample complexity of learning DR-DMDPs has been extensively studied, the average-reward setting remains relatively underexplored. Wang et al., 2023e ; Wang et al., 2023d ; Wang et al., 2024d propose Robust relative value iteration and TD/Q-learning algorithms and prove their convergence, but without providing sample complexity guarantees. Grand-Clément et al., (2025) show that for $(s,a)$-rectangular uncertainty sets, the optimal policy can be stationary and deterministic; however, this result may not extend to $s$-rectangular uncertainty sets. Although existing work has investigated the existence and structure of optimal policies in average-reward DR-RL, non-asymptotic sample complexity bounds remain an open question.
@@
Standard
Discounted $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-3}\epsilon^{-2}})$ Azar et al., (2013); Li et al., (2020)
Discounted Mixing $\widetilde{\Theta}\left(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}(1-\gamma)^{-2}\varepsilon^{-2}\right)$ Wang et al., 2023a
Average Mixing $\widetilde{\Theta}\left(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}\varepsilon^{-2}\right)$ Wang et al., 2024a ; Zurek and Chen, (2023)
DR-RL
Discounted $\widetilde{O}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-4}\epsilon^{-2}})$ Shi and Chi, (2024); Wang et al., 2024c
@@ 3.1 Markov Decision Processes
Define the canonical space $\Omega=(\mathbf{S}\times\mathbf{A})^{\mathbb{N}}$ equipped with $\mathcal{F}$ the $\sigma$-field generated by cylinder sets. The state-action process $\left\{{(S_{t},A_{t}),{t\geq 0}}\right\}$ is defined by the point evaluation $X_{t}(\omega)=s_{t},A_{t}(\omega)=a_{t}$ for all $t\geq 0$ for any $\omega=(s_{0},a_{0},s_{1},a_{1},\dots)\in\Omega$. A general history dependent policy $\pi=(\pi_{t})_{t\geq 0}\in\Pi_{\mathrm{HD}}$ is a sequence of the agent’s decision rule. Here, the decision rule $\pi_{t}$ at time $t$ is a mapping $\pi_{t}:(\mathbf{S}\times\mathbf{A})^{t}\times\mathbf{S}\to\Delta(\mathbf{A})$, signifying the conditional distribution of $A_{t}$ given the history. It is known in the literature (Puterman,, 2009; Grand-Clément et al.,, 2025) that to achieve optimal decision making in the context of infinite horizon AMDPs, DMDPs, or their robust variants (to be introduced), it suffices to consider the policy class $\Pi$ of stationary, Markov, and deterministic policies; i.e. $\pi\in\Pi$ can be seen as a function $\pi:\mathbf{S}\rightarrow\mathbf{A}$. Thus, in the subsequent development, we restrict our discussion to $\Pi$.
As in Wang et al., 2024a a policy $\pi\in\Pi$ and an initial distribution $\mu\in\Delta(\mathbf{S})$ uniquely defines a probability measure on $(\Omega,\mathcal{F})$. We will always assume that $\mu$ is the uniform distribution over $\mathbf{S}$. The expectation under this measure is denoted by $E_{P}^{\pi}$. To simplify notation, we define $P_{\pi}(s,s^{\prime}):=\sum_{a\in\mathbf{A}}\pi(a|s)p_{s,a}(s^{\prime})$ and $r_{\pi}(s):=\sum_{a\in\mathbf{A}}\pi(a|s)r(s,a)$.
@@
When $P$ is uniformly ergodic (a.k.a.unichain, to be defined later), the $g_{P}^{\pi}$ is contstant across states (Puterman,, 2009). In this context, an optimal policy $\pi^{*}\in\Pi$ achieves the long-run average-reward $\max_{\pi\in\Pi}g_{P}^{\pi}$.
@@ 3.2 Uniform Ergodicity
It is well known (Meyn and Tweedie,, 2012) that $\rho$ must be the unique stationary distribution of $K$ and that two conditions are equivalent. The $\psi$ and $R$ in the Doeblin condition are known as the minorization measure and the residual kernel respectively.
@@
It is shown in Theorem 1 of Wang et al., 2023a that for a uniformly ergodic transition kernel $K$, these metrics of stability are equivalent up to constants: $t_{\mathrm{minorize}}(K)\leq 22t_{\mathrm{mix}}(K)\leq 22\log(16)t_{\mathrm{minorize}}(K)$.
@@ 3.3 Distributionally Robust Discounted-Reward and Average-Reward MDPs
For each $(s,a)\in\mathbf{S}\times\mathbf{A}$ we define the uncertainty set under divergence $D=D_{\mathrm{KL}},D_{f_{k}}$ and parameter $\delta>0$ centered at $p_{s,a}$ by $\mathcal{P}_{s,a}(D,\delta):=\left\{{p:D(p\|p_{s,a})\leq\delta}\right\}$. This set contains all possible adversarial perturbations of the transition out of $(s,a)$. Note that the parameter $\delta$ controls the size of the $\mathcal{P}_{s,a}(D,\delta)$, quantifying the power of the adversary. The uncertainty set for the entire controlled transition kernel is $\mathcal{P}(D,\delta):=\bigtimes_{(s,a)\in\mathbf{S}\times\mathbf{A}}\mathcal{P}_{s,a}(D,\delta)$. An uncertainty set of this product from is called SA-rectangular (Wang et al., 2024b, ).
@@
See Iyengar, (2005) for a rigorous construction of the expectation $E_{\mathbf{P}}^{\pi}$. Then, the optimal value function is $V_{\mathcal{P}}^{*}(s):=\max_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}(s)$. It is well known (c.f. Iyengar, (2005)) that $V_{\mathcal{P}}^{*}$ is the unique solution of the DR Bellman equation: $V_{\mathcal{P}}^{*}=\mathcal{T}_{\gamma}^{*}(V_{\mathcal{P}}^{*})$.
Note that the expectation $E_{\mathbf{P}}^{\pi}$ is under the adversarial perturbation from a Markovian policy class $(\mathcal{P})^{\mathbb{N}}$. It is possible to consider other information structures for the adversary while retaining the satisfaction of the Bellman equation (Wang et al., 2024b, ).
@@
We note that $\mathcal{P}=\mathcal{P}(D,\delta)$ is compact in the sense that $\mathcal{P}_{s,a}(D,\delta)$ is a compact subset of $\Delta(\mathbf{S})$ for all $s,a$. With uniform ergodicity and compactness, Wang et al., 2023d shows that $g_{\mathcal{P}}^{*}(s)$ is constant for $s\in\mathbf{S}$ which uniquely solves the DR Bellman equations.
Proposition 3.6 (Theorem 7 of Wang et al., 2023d ).
If $\mathcal{P}$ is uniformly ergodic, then $g_{\mathcal{P}}^{*}(s)\equiv g_{\mathcal{P}}^{*}$ is constant in $s\in\mathbf{S}$. Moreover, a solution $(g,v)$ of $v(s)=\mathcal{T}^{*}(v)(s)-g^{*}(s)$ for all $s\in\mathbf{S}$ satisfies $g(s)=g_{\mathcal{P}}^{*}$ for all $s,s^{\prime}\in\mathbf{S}$. Moreover, the policy $\pi^{*}(s)\in\arg\max_{a\in\mathbf{A}}\left\{{r(s,a)+\Gamma_{\mathcal{P}_{s,a}}(v)}\right\}$ achieves the optimal average-reward $g_{\mathcal{P}}^{*}$.

### v3

@@ 1 Introduction
Reinforcement learning (RL) [35] is a core machine learning framework in which agents learn to make decisions by interacting with their environments to maximize long-term rewards. RL has been successfully applied across a wide range of domains—from classic applications in robotics and control systems [19, 13] to more recent advances in game playing [21, 6, 8] and large language model (LLM)-driven reasoning tasks [47, 14].
@@
To address this challenge, Zhou et al., [56] built upon the distributionally robust Markov decision process (DR-MDP) framework [16, 27, 48] to propose a distributionally robust reinforcement learning (DR-RL) framework. Subsequent work advanced the field, including both model-free [22, 40, 43] and model-based settings [29, 50, 7, 33], as well as approaches for offline learning [32] and generative models [40, 43, 53, 7], along with functional approximations [2, 24].
@@
Average-reward RL is not only important but also more challenging in terms of algorithm design and theoretical analysis. In the standard (non-robust) RL setting, the minimax sample complexity for generative models in discounted-reward cases was resolved as early as 2013 [1]. In contrast, analogous results for average-reward settings were only developed much later, with recent advances in Wang et al., 2023a [37], Zurek and Chen, [57], Zurek and Chen, 2024a [58], Zurek and Chen, 2024b [59], Zurek and Chen, [60] under granular structural assumptions.
@@ 2 Literature Review
Standard Discounted $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-3}\epsilon^{-2}})$ Azar et al., [1], Li et al., [20]
Discounted Mixing $\widetilde{\Theta}\left(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}(1-\gamma)^{-2}\varepsilon^{-2}\right)$ Wang et al., 2023a [37]
Average Mixing $\widetilde{\Theta}\left(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}\varepsilon^{-2}\right)$ Wang et al., 2024a [38], Zurek and Chen, [57]
DR-RL Discounted $\widetilde{O}(|\mathbf{S}||\mathbf{A}|{(1-\gamma)^{-4}\epsilon^{-2}})$ Shi and Chi, [32], Wang et al., 2024c [43]
@@
Sample Complexity of Discounted-Reward Tabular RL: There is an extensive literature on the sample complexity of tabular reinforcement learning. In standard (non-robust) settings, the minimax sample complexity for discounted-reward problems has been well studied. Early works by Azar et al., [1], Li et al., [20] established the minimax rate of $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|(1-\gamma)^{-3}\varepsilon^{-2})$. More recent research has shifted toward instance-dependent bounds that leverage structural properties of the MDP. For example, Wang et al., 2024a [38], Zurek and Chen, 2024b [59] derive tighter instance-dependent bounds of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}(1-\gamma)^{-2}\varepsilon^{-2})$ and $\widetilde{O}(|\mathbf{S}||\mathbf{A}|\mathrm{H}(1-\gamma)^{-2}\varepsilon^{-2})$ under the assumptions that $P$ is uniformly ergodic or weakly communicating, where $\mathrm{H}$ denotes the span of the relative value function.
Sample Complexity of Average-Reward Tabular RL: Recently, there has been growing interest in the sample complexity of average-reward reinforcement learning in standard (non-robust) settings. Early work by Jin and Sidford, [17] established a bound of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}^{2}\varepsilon^{-2})$ using primal-dual stochastic mirror descent. A rate of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}\varepsilon^{-3})$ was established later via a reduction to the discounted MDP setting [18]. Subsequent analyses achieved tighter bounds: Wang et al., 2024a [38] obtained a rate of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}\varepsilon^{-2})$, while Zurek and Chen, [57], Zurek and Chen, 2024b [59] derived instance-dependent optimal rates of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|\mathrm{H}\varepsilon^{-2})$ for weakly communicating MDPs. Further refinements [58, 60] achieved similar rates without requiring prior knowledge, using plug-in and span penalization approaches. Beyond the generative model setting, Zhang et al., [55] and Chen, [5] provided finite-sample analyses for synchronous and asynchronous Q-learning, respectively. Asymptotic properties were also studied in Yu et al., [54], Wan et al., [36] for asynchronous $Q$-learning.
DR-DMDP and DR-RL: Our work builds on the theoretical foundations of robust MDPs [11, 16, 27, 48, 49, 31, 42], which primarily develop dynamic programming principles under the discounted-reward setting. Recent advances in distributionally robust reinforcement learning (DR-RL) have investigated the sample complexity of DR-DMDPs under various divergence-based uncertainty sets. For example, Wang et al., 2024c [43], Shi and Chi, [32] establish a model-free upper bound of $\widetilde{O}(|\mathbf{S}||\mathbf{A}|(1-\gamma)^{-4}\varepsilon^{-2})$ under KL-divergence. Under $\chi^{2}$-divergence, Shi et al., [33] obtain a similar upper bound, while Clavier et al., [7] show that $l_{p}$-norm constraints admit a tighter minimax rate of $\widetilde{\Theta}(|\mathbf{S}||\mathbf{A}|(1-\gamma)^{-3}\varepsilon^{-2})$. In this paper, by incorporating the mixing time parameter, we present a “instance-dependence" sample complexity bound of $\widetilde{O}\left(|\mathbf{S}||\mathbf{A}|t_{\mathrm{mix}}^{2}(1-\gamma)^{-2}\varepsilon^{-2}\right)$, which improves the dependence on the effective horizon from $(1-\gamma)^{-4}$ to $(1-\gamma)^{-2}$. Several other works contribute to the theoretical and algorithmic landscape of DR-RL in various settings, including Panaganti and Kalathil, [28], Yang et al., [52], Xu et al., [50], Blanchet et al., [3], Liu et al., 2022b [23], Wang et al., 2023c [41], Yang et al., [51].
Distributionally Robust Average-Reward MDPs: While the sample complexity of learning DR-DMDPs has been extensively studied, the average-reward setting remains relatively underexplored. Wang et al., 2023e [46], Wang et al., 2023d [44], Wang et al., 2024d [45] propose robust relative value iteration and TD/Q-learning algorithms and prove their convergence, but without providing sample complexity guarantees. Grand-Clément et al., [12] show that for $(s,a)$-rectangular uncertainty sets, the optimal policy can be stationary and deterministic; however, this result may not extend to $s$-rectangular uncertainty sets. More recently, Wang and Si, [39] study average-reward robust MDPs under $s$-rectangular uncertainty and provide one-sided weak communication conditions that ensure the existence of solutions to the Bellman optimality equation. Although existing work has investigated the existence and structure of optimal policies in average-reward DR-RL, non-asymptotic sample complexity bounds remain an open question.
@@ 3.1 Markov Decision Processes
Define the canonical space $\Omega=(\mathbf{S}\times\mathbf{A})^{\mathbb{N}}$ equipped with $\mathcal{F}$ the $\sigma$-field generated by cylinder sets. The state-action process $\left\{{(S_{t},A_{t}),{t\geq 0}}\right\}$ is defined by the point evaluation $X_{t}(\omega)=s_{t},A_{t}(\omega)=a_{t}$ for all $t\geq 0$ for any $\omega=(s_{0},a_{0},s_{1},a_{1},\dots)\in\Omega$. A general history dependent policy $\pi=(\pi_{t})_{t\geq 0}\in\Pi_{\mathrm{HD}}$ is a sequence of the agent’s decision rule. Here, the decision rule $\pi_{t}$ at time $t$ is a mapping $\pi_{t}:(\mathbf{S}\times\mathbf{A})^{t}\times\mathbf{S}\to\Delta(\mathbf{A})$, signifying the conditional distribution of $A_{t}$ given the history. It is known in the literature [30, 12] that to achieve optimal decision making in the context of infinite horizon AMDPs, DMDPs, or their robust variants (to be introduced), it suffices to consider the policy class $\Pi$ of stationary, Markov, and deterministic policies; i.e. $\pi\in\Pi$ can be seen as a function $\pi:\mathbf{S}\rightarrow\mathbf{A}$. Thus, in the subsequent development, we restrict our discussion to $\Pi$.
As in Wang et al., 2024a [38] a policy $\pi\in\Pi$ and an initial distribution $\mu\in\Delta(\mathbf{S})$ uniquely defines a probability measure on $(\Omega,\mathcal{F})$. We will always assume that $\mu$ is the uniform distribution over $\mathbf{S}$. The expectation under this measure is denoted by $E_{P}^{\pi}$. To simplify notation, we define $P_{\pi}(s,s^{\prime}):=\sum_{a\in\mathbf{A}}\pi(a|s)p_{s,a}(s^{\prime})$ and $r_{\pi}(s):=\sum_{a\in\mathbf{A}}\pi(a|s)r(s,a)$.
@@
When $P$ is uniformly ergodic (a.k.a.unichain, to be defined later), the $g_{P}^{\pi}$ is constant across states [30]. In this context, an optimal policy $\pi^{*}\in\Pi$ achieves the long-run average-reward $\max_{\pi\in\Pi}g_{P}^{\pi}$.
@@ 3.2 Uniform Ergodicity
It is well known [25] that $\rho$ must be the unique stationary distribution of $K$ and that two conditions are equivalent. The $\psi$ and $R$ in the Doeblin condition are known as the minorization measure and the residual kernel respectively.
@@
It is shown in Theorem 1 of Wang et al., 2023a [37] that for a uniformly ergodic transition kernel $K$, these metrics of stability are equivalent up to constants: $t_{\mathrm{minorize}}(K)\leq 22t_{\mathrm{mix}}(K)\leq 22\log(16)t_{\mathrm{minorize}}(K)$.
@@ 3.3 Distributionally Robust Discounted-Reward and Average-Reward MDPs
For each $(s,a)\in\mathbf{S}\times\mathbf{A}$ we define the uncertainty set under divergence $D=D_{\mathrm{KL}},D_{f_{k}}$ and parameter $\delta>0$ centered at $p_{s,a}$ by $\mathcal{P}_{s,a}(D,\delta):=\left\{{p:D(p\|p_{s,a})\leq\delta}\right\}$. This set contains all possible adversarial perturbations of the transition out of $(s,a)$. Note that the parameter $\delta$ controls the size of the $\mathcal{P}_{s,a}(D,\delta)$, quantifying the power of the adversary. The uncertainty set for the entire controlled transition kernel is $\mathcal{P}(D,\delta):=\bigtimes_{(s,a)\in\mathbf{S}\times\mathbf{A}}\mathcal{P}_{s,a}(D,\delta)$. An uncertainty set of this product from is called SA-rectangular [42].
@@
See Iyengar, [16] for a rigorous construction of the expectation $E_{\mathbf{P}}^{\pi}$. Then, the optimal value function is $V_{\mathcal{P}}^{*}(s):=\max_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}(s)$. It is well known (c.f. Iyengar, [16]) that $V_{\mathcal{P}}^{*}$ is the unique solution of the DR Bellnbman equation: $V_{\mathcal{P}}^{*}=\mathcal{T}_{\gamma}^{*}(V_{\mathcal{P}}^{*})$.
Note that the expectation $E_{\mathbf{P}}^{\pi}$ is under the adversarial perturbation from a Markovian policy class $(\mathcal{P})^{\mathbb{N}}$. It is possible to consider other information structures for the adversary while retaining the satisfaction of the Bellman equation [42].
@@
We note that $\mathcal{P}=\mathcal{P}(D,\delta)$ is compact in the sense that $\mathcal{P}_{s,a}(D,\delta)$ is a compact subset of $\Delta(\mathbf{S})$ for all $s,a$. With uniform ergodicity and compactness, Wang et al., 2023d [44] shows that $g_{\mathcal{P}}^{*}(s)$ is constant for $s\in\mathbf{S}$ which uniquely solves the DR Bellman equations.
Proposition 3.6 (Theorems 7 an 8 of Wang et al., 2023d [44]).
If $\mathcal{P}$ is uniformly ergodic with a uniformly bounded minorization time, then $g_{\mathcal{P}}^{*}(s)\equiv g_{\mathcal{P}}^{*}$ is constant in $s\in\mathbf{S}$. Moreover, there exists a solution $(g,v)$ of $v(s)=\mathcal{T}^{*}(v)(s)-g^{*}(s)$ for all $s\in\mathbf{S}$ and any such solution satisfies $g(s)=g_{\mathcal{P}}^{*}$ for all $s\in\mathbf{S}$. Moreover, the policy $\pi^{*}(s)\in\arg\max_{a\in\mathbf{A}}\left\{{r(s,a)+\Gamma_{\mathcal{P}_{s,a}}(v)}\right\}$ achieves the optimal average-reward $g_{\mathcal{P}}^{*}$.